# Universal Mott quantum criticality in a modified periodic Anderson model

### v1

@@ III.2 $\omega/T$ scaling at the QCP:
Figure 3: (top)Imaginary part of the self-energy, $\Sigma_{\text{f}}(\omega)$ and (bottom)spectral function, $\text{A}_{\text{f}}(\omega)$, are plotted for different $\beta$’s at the QCP. We can observe $\omega/T$ scaling in both of them with exponents $0.36,-0.36(\text{A}_{\text{f}}(\omega),\text{Im}\Sigma_{\text{f}}(\omega))$. (Inset) shows the plot of unscaled $A(\omega)$ vs. $\omega$. Parameters used: $U=1.75$.
Figure 4: Scaling plot of the dynamical susceptibility, $T^{0.36}\chi^{\prime\prime}(\omega,T)$, shown for several temperatures ($\beta$ ranging from 50 to 1000). (a) At the quantum critical point ($t_{\perp}=t_{\perp,c}=1.033$), the data collapse onto a universal curve over three decades, consistent with quantum critical scaling [32]. (b, c) Away from the QCP, for $t_{\perp}=0.7$ (top right) and $t_{\perp}=1.1$ (bottom right), the scaling collapse clearly fails. Parameters used: $U=1.75$.
The SBHM[59], Kondo lattice model[60], pseudogap Anderson model[32], and the pseudogap Kondo model display “local” quantum criticality and are often associated with the $\omega/T$ scaling in the quantum critical region. Numerous neutron scattering experiments have reported quantum criticality manifested through $\omega/T$ scaling [61, 62, 63, 64], which belong to the class of “local” quantum criticality[65, 66, 60, 67]. In this section, we investigate the presence of such $\omega/T$ scaling behavior in the MPAM.
As a first step towards identifying such signatures, we analyze the single-particle dynamics [32]. Fig. 3 shows $\omega/T$ scaling in both the scaled spectral function, $T^{-r}\text{A}_{\text{f}}(\omega)$, and the scaled imaginary part of the self-energy, $T^{r}\,\text{Im}\,\Sigma_{\text{f}}(\omega)$. The scaling exponent is found to be $r=0.36$, consistent with the value reported in Ref. 28. The inset in the bottom part of Fig. 3 displays the spectral function for various temperatures at the QCP.
@@
where $S^{z}=\frac{1}{2}(n_{\uparrow}-n_{\downarrow})$ is the local spin operator. To this end, we examine the two-particle response by computing the scaled imaginary part of the dynamical spin-spin susceptibility, $T^{r}\chi^{\prime\prime}(\omega,T)(r=0.36)$, where $\chi^{\prime\prime}(\omega,T)$ denotes the imaginary part of $\chi(\omega,T)$. At the QCP ($t_{\perp}=1.033$), we observe clear $\omega/T$ scaling in $T^{0.36}\chi^{\prime\prime}(\omega,T)$, as shown in Fig. 4. Importantly, such scaling is observed only at the critical point; as the system is tuned away from the QCP, the scaling behavior disappears as shown in the Fig. 4.
To conclude this subsection, we establish that the MPAM exhibits local quantum criticality by demonstrating $\omega/T$ scaling in both single-particle and two-particle correlation functions at the QCP. In the following, we discuss the transport properties of the MPAM.
@@ III.3 Transport quantities
In this section, we analyze the optical conductivity to identify signatures of the quantum critical point. We further compute the DC conductivity and resistivity, demonstrating a scaling collapse of the latter with an exponent comparable to that of the SBHM.
@@ III.3.1 Optical conductivity
Figure 5: (top) Optical conductivity plotted for various inverse temperatures $\beta$ at $t_{\perp}=0.7$. The inset shows two isosbestic points located before and after the MIR peak. (bottom) Optical conductivity plotted for $t_{\perp}=1.033$ (top) and $1.08$ (bottom). At the QCP, the conductivity $\sigma(\omega)$ follows a power-law behavior, as indicated by the dashed maroon line, with an exponent of $0.737$, which is approximately $2r$, where $r=0.36$ is the critical exponent[28]. Parameters used: $U=1.75$.
A general expression for the optical conductivity, based on the Kubo formula, is given by $\sigma(\omega)=\frac{1}{i\omega^{+}}\,\langle\langle j;j\rangle\rangle_{\omega},$ where $\langle\langle j;j\rangle\rangle_{\omega}$ denotes the retarded current-current correlation function. Within DMFT, the self-energy is purely local, and vertex corrections vanish in the skeleton expansion of the current–current correlation function. Consequently, only the bare bubble diagram contributes to the conductivity, which implies that single-particle quantities are sufficient to compute $\sigma(\omega)$. The optical conductivity is then calculated using the standard expression [19, 23, 25, 24]:
@@
where $\mathbf{A}(\omega,\epsilon)$ is a $2\times 2$ spectral function matrix with diagonal elements -$G_{\text{cc}}(\omega,\epsilon), -Im$$G_{\text{mm}}(\omega,\epsilon)$, and off-diagonal elements -$G_{\text{mc}}(\omega,\epsilon), -Im$$G_{\text{cm}}(\omega,\epsilon)$. $\text{A}_{0}$ is the non-interacting Bethe lattice density of states[25, 68, 24], $v(\epsilon)=\sqrt{D^{2}-\epsilon^{2}}/\sqrt{3}$ represents the energy-dependent velocity on the Bethe lattice, $n_{F}(\omega)=(1+e^{\beta\omega})^{-1}$ is the Fermi–Dirac distribution, and $\sigma_{0}$ is a material-dependent constant. This expression holds in the $z\rightarrow\infty$ limit of the Bethe lattice.
@@
The optical conductivity is computed for various values of inter-orbital coupling $t_{\perp}$ and inverse temperature $\beta$, as shown in Fig. 5. For $t_{\perp}=0.7$, we observe three distinct features: (1) a low-frequency Drude peak, originating from coherent quasiparticle excitations near the Fermi level; (2) a mid-infrared peak arising from transitions between the lower Hubbard band and the quasiparticle resonance, and (3) a high-energy charge-excitation peak corresponding to transitions between the lower and upper Hubbard bands. With decreasing temperature, the Drude peak becomes narrower and its height increases, whereas the charge-excitation peak remains largely temperature independent, similar to the earlier reports in the doped infinite-dimensional Hubbard model [69]. As $t_{\perp}$ approaches the critical value $t_{\perp,c}$, the mid-IR peak is suppressed. For $t_{\perp}>t_{\perp,c}$, an optical gap develops due to the absence of coherent quasiparticles.
@@
In the metallic regime, the optical conductivity spectrum for $\omega\geq 0$ is composed of three distinct features. The first is a Drude peak at zero frequency. The other two features arise from three possible intraband transitions: $\Delta_{ul}$, $\Delta_{um}$, and $\Delta_{ml}$. Of these, the $\Delta_{um}$ and $\Delta_{ml}$ transitions are degenerate, combining to form a single mid-infrared (MIR) peak. The remaining transition, $\Delta_{ul}$, gives rise to the charge excitation peak. Consequently, the total optical spectrum clearly exhibits these three peaks. The charge-excitation peak in $\sigma(\omega)$, corresponding to the transition from the lower to the upper band, is minimum when $\epsilon=0$, and is approximately given by $\Delta_{ul}(\epsilon=0)=t_{\perp}+\sqrt{t_{\perp}^{2}+2ZV^{2}}\approx 2t_{\perp}$. It is interesting to note that, in the limit $t_{\perp}=0$, the MPAM reduces to the PAM, and $\Delta_{ul}\approx\sqrt{2ZV^{2}}$, which is smaller by a factor of $\sqrt{2}$ compared to the gap obtained for the PAM [70] ($2\sqrt{ZV^{2}}$). This discrepancy arises from the approximation used in determining the roots of the MPAM (see Eq. 11), where the term $\frac{ZV^{2}}{2\omega t_{\perp}}$ is neglected. Including this term would yield the same result as the PAM in the limit $t_{\perp}\rightarrow 0$.
@@
The MIR peak originates from either $\Delta_{um}$ or $\Delta_{ml}$. Their contributions are identical due to particle–hole symmetry, with the minimum excitation occurring at $\epsilon=-\min(D,t_{\perp})$ for $\Delta_{um}$ and at $\epsilon=\min(D,t_{\perp})$ for $\Delta_{ml}$. Hence, the position of the MIR peak, $\omega_{\text{MIR}}$, is proportional to $\Delta_{um}(\epsilon=-\text{min}(D,t_{\perp}))=\Delta_{ml}(\epsilon=\text{min}(D,t_{\perp}))\approx\sqrt{|\text{min}(D,t_{\perp})-t_{\perp}|^{2}+2ZV^{2}}$. As the system approaches the critical point, both $Z\rightarrow 0$ and $|\text{min}(D,t_{\perp})-t_{\perp}|$ decrease. Consequently, the MIR peak redshifts towards $\omega=0$ and the charge-excitation peak blue-shifts away from $\omega=0$ as the system approaches the QCP. In particular, $\omega_{\text{MIR}}$ reaches its minimum at the QCP with an exponent of $0.62$, as shown in the top panel of Fig.6. Notably, the MIR peak persist even in the $Z\rightarrow 0$ limit due to the finite contribution from $|D-t_{\perp,c}|$ ($t_{\perp,c}=1.033>D$). This trend is evident in Fig. 7, where the MIR peak moves towards $\omega=0$ as $t_{\perp}$ increases, reaching a minimum at $t_{\perp}=t_{\perp,c}$. For $t_{\perp}<D=1.0$, however, $\omega_{\text{MIR}}\propto\sqrt{2ZV^{2}}$, which is explicitly independent of $t_{\perp}$. This implies that the MIR peak provides information about the hybridization gap and quasi-particle weight, similar to the case in the PAM [57]. However, as $t_{\perp}\rightarrow t_{\perp,c}$, other model parameters (here, $t_{\perp}$) can also influence the MIR peak.
Beyond the QCP ($t_{\perp}>t_{\perp c}$), the MIR peak disappears. At the QCP, due to the diverging nature of $\Sigma(\omega)$, the spectral functions, $A(\omega,\epsilon)\propto\omega^{r}$, with r=0.36r=0.36[28] being the soft-gap exponent. This results in a power-law behavior of the optical conductivity with an exponent $2r$. The bottom panel of Fig. 5 shows that as the temperature is decreased, the optical conductivity approaches a power-law form with an exponent $0.737$, which is approximately $2r$.
@@
In the regime $t_{\perp}<t_{\perp,c}$, two ISPs can be observed, one before the MIR peak ($\omega_{1}^{\text{ISP}}<\omega_{\text{MIR}}$) and another after it ($\omega_{2}^{\text{ISP}}>\omega_{\text{MIR}}$). These ISPs gradually disappear as the system approaches the QCP. We track them by extracting the corresponding frequencies, $\omega_{1/2}^{\text{ISP}}$. The bottom panel of Fig. 6 presents the evolution of ISPs as a function of $t_{\perp}$. We observe that $\omega_{1}^{\text{ISP}}$ vanishes linearly at the QCP, while $\omega_{2}^{\text{ISP}}$ disappears before reaching the QCP. Observing these distinct behaviors experimentally would offer compelling evidence for the existence of a QCP.
We have investigated the optical conductivity and identified clear signatures of a QCP. Next, we calculate the DC conductivity and resistivity and analyze the quantum critical scaling behavior of the resistivity.
@@ III.3.2 DC conductivity and resistivity scaling
Figure 8: The DC conductivity, $\sigma^{\text{DC}}(T)$, computed using Eq. 9, is plotted for several values of the inter-orbital coupling $t_{\perp}$. The blue dashed line denotes the result from a toy model calculation at the QCP. As $T\rightarrow 0$ at the QCP, the toy model DC conductivity, $\sigma^{\text{DC}}{\text{QCP}}$, matches $\sigma^{\text{DC}}(T,t{\perp,c})$. Parameters: $U=1.75$.
The DC conductivity is determined by taking the limit as $\omega\rightarrow 0$ in the optical conductivity expression (7), which then reduces to the following form:
$\frac{\sigma^{\text{DC}}(T)}{\sigma_{0}}$ $=2\pi\int d\epsilon\,\text{A}_{0}(\epsilon)\int d\omega^{\prime}\left(-\frac{dn_{F}}{d\omega^{\prime}}\right)$ (8)
@@
DC conductivity, $\sigma^{\text{DC}}(T)$, is plotted for different inter-orbital couplings, $t_{\perp}$, in Fig. 7. As expected, $\sigma^{\text{DC}}(T)$ diverges for $t_{\perp}<t_{\perp,c}$ and vanishes for $t_{\perp}>t_{\perp,c}$. At the QCP, it follows a different power-law behavior. At the QCP, the characteristic $\omega/T$ scaling implies a power-law form for the self-energy, $\Sigma_{\text{f}}(\omega)\propto T^{-r}|\omega/T|^{-r}$. This, in turn, governs the behavior of the spectral functions, resulting in the relation $\text{A}_{\text{c}}(\omega,\epsilon)=\text{A}_{\text{c}_{M}}(\omega,\epsilon)\propto T^{2r}$. This leads to the DC conductivity scaling as $\sigma_{\text{DC}}\propto T^{2r}$ at the QCP. A toy model calculation suggests that very low temperatures, as low as $T=10^{-5}$, are required to observe this scaling clearly. However, such calculations suffer from oscillations at very low frequencies and temperatures [68]. Ideally, accurate results would be obtained in the limit $\Lambda\rightarrow 1$ with a larger $n_{\text{states}}$, but this remains infeasible with current computational resources.
Figure 9: (Main)Scaled DC resistivity, $\rho^{DC}(T)/\rho^{DC}_{c}(T)$ is plotted as a function of scaled temperature($T/T_{0}$) for different values of $U$. (Inset) $T_{0}$ is plotted as a function of $\delta U=|U-U_{C}|$, where $U_{c}=1.75$. $T_{0,\text{M}},T_{0,\text{I}}$ corresponds to the scaled temperatures $T_{0}$ in Metal and Mott insulating side. $\text{T}_{0,\text{M}},T_{0,\text{I}}$ vanishes with exponents ($0.66,0.76$). Parameters used: $t_{\perp}=t_{\perp\text{c}}=1.033$.
Resistivity is obtained from the inverse of the DC conductivity, $\rho^{\text{DC}}=1/\sigma^{\text{DC}}$. Systems exhibiting Mott quantum criticality display a characteristic scaling behavior in the electrical resistivity [23, 25, 24]. This scaling provides insights into the quantum critical region and allows for the extraction of the critical exponent $\nu z$, where $\nu$ is the correlation length exponent and $z$ is the dynamical exponent. A hallmark of Mott quantum criticality is the scaling collapse of the DC resistivity curves. This universal behavior is captured by the scaling form:

### v2

@@ III.2 $\omega/T$ scaling at the QCP:
Figure 3: (top)Imaginary part of the self-energy, $\Sigma_{\text{f}}(\omega)$ and (bottom)spectral function, $\text{A}_{\text{f}}(\omega)$, are plotted for different $\beta$’s at the QCP. We can observe $\omega/T$ scaling in both of them with exponents $0.36\,(\text{A}_{\text{f}}(\omega)),-0.36\,(\text{Im}\Sigma_{\text{f}}(\omega))$. (Inset) shows the plot of unscaled $A(\omega)$ vs. $\omega$. Parameters used: $U=1.75$.
Figure 4: Scaling plot of the dynamical susceptibility, $T^{0.36}\chi^{\prime\prime}(\omega,T)$, shown for several temperatures ($\beta$ ranging from 50 to 1000). (a) At the quantum critical point ($t_{\perp}=t_{\perp,c}=1.033$), the data collapse onto a universal curve over three decades, consistent with quantum critical scaling [glossop2011critical]. (b, c) Away from the QCP, for $t_{\perp}=0.7$ (top right) and $t_{\perp}=1.1$ (bottom right), the scaling collapse clearly fails. Parameters used: $U=1.75$.
The SBHM [dasari2017quantum], Kondo lattice model [si2003local], pseudogap Anderson model [glossop2011critical], and the pseudogap Kondo model display “local” quantum criticality and are often associated with the $\omega/T$ scaling in the quantum critical region. Numerous neutron scattering experiments have reported quantum criticality manifested through $\omega/T$ scaling  [aronson1995non, stockert1998two, schroder2000onset, michon2023reconciling], which belong to the class of “local” quantum criticality [si2001locally, ingersent2002critical, si2003local, gegenwart2008quantum]. In this section, we investigate the presence of such $\omega/T$ scaling behavior in the MPAM.
As a first step towards identifying such signatures, we analyze the single-particle dynamics [glossop2011critical]. Fig. 3 shows $\omega/T$ scaling in both the scaled spectral function, $T^{-r}\text{A}_{\text{f}}(\omega)$, and the scaled imaginary part of the self-energy, $T^{r}\,\text{Im}\,\Sigma_{\text{f}}(\omega)$. The scaling exponent is found to be $r=0.36$, consistent with the value reported in Ref. kksujan2023emergent. The inset in the bottom part of Fig. 3 displays the spectral function for various temperatures at the QCP.
@@
where $S^{z}=\frac{1}{2}(n_{\uparrow}-n_{\downarrow})$ is the local spin operator. To this end, we examine the two-particle response by computing the scaled imaginary part of the dynamical spin-spin susceptibility, $T^{r}\chi^{\prime\prime}(\omega,T)(r=0.36)$, where $\chi^{\prime\prime}(\omega,T)$ denotes the imaginary part of $\chi(\omega,T)$. At the QCP ($t_{\perp}=1.033$), we observe clear $\omega/T$ scaling in $T^{0.36}\chi^{\prime\prime}(\omega,T)$, as shown in Fig. 4. Importantly, such scaling is observed only at the critical point; as the system is tuned away from the QCP, the scaling behavior disappears as shown in the Fig.  4.
The demonstration of robust $\omega/T$ scaling in both single- and two-particle correlation functions establishes that the QCP in the MPAM is locally critical. Building on this finding, we now turn to the transport properties of the MPAM to identify the signatures of Mott quantum criticality.
@@ III.3 Transport quantities
In this section, we analyze the optical conductivity to identify signatures of the Mott quantum criticality. We further compute the DC conductivity and resistivity, demonstrating a scaling collapse of the latter with an exponent comparable to that of the SBHM.
@@ III.3.1 Optical conductivity
Figure 5: (top) Optical conductivity plotted for various inverse temperatures $\beta$ at $t_{\perp}=0.7$. The inset shows two isosbestic points located before and after the MIR peak. (bottom) Optical conductivity plotted for $t_{\perp}=1.033$ (top) and $1.08$ (bottom). At the QCP, the conductivity $\sigma(\omega)$ follows a power-law behavior, as indicated by the dashed maroon line, with an exponent of $0.737$, which is approximately $2r$, where $r=0.36$ is the critical exponent [kksujan2023emergent]. Parameters used: $U=1.75$.
A general expression for the optical conductivity, based on the Kubo formula, is given by $\sigma(\omega)=\frac{1}{i\omega^{+}}\,\langle\langle j;j\rangle\rangle_{\omega},$ where $\langle\langle j;j\rangle\rangle_{\omega}$ denotes the retarded current-current correlation function. Within DMFT, the self-energy is purely local, and vertex corrections vanish in the skeleton expansion of the current–current correlation function. Consequently, only the bare bubble diagram contributes to the conductivity, which implies that single-particle quantities are sufficient to compute $\sigma(\omega)$ [georges1996dynamical]. The optical conductivity is then calculated using the standard expression [georges1996dynamical, terletska2011quantum, vuvcivcevic2013finite, eisenlohr2019mott]:
@@
where $\mathbf{A}(\omega,\epsilon)$ is a $2\times 2$ spectral function matrix with diagonal elements -$G_{\text{cc}}(\omega,\epsilon), -Im$$G_{\text{mm}}(\omega,\epsilon)$, and off-diagonal elements -$G_{\text{mc}}(\omega,\epsilon), -Im$$G_{\text{cm}}(\omega,\epsilon)$. $\text{A}_{0}$ is the non-interacting Bethe lattice density of states [vuvcivcevic2013finite, vzitko2015repulsive, eisenlohr2019mott], $v(\epsilon)=\sqrt{D^{2}-\epsilon^{2}}/\sqrt{3}$ represents the energy-dependent velocity on the Bethe lattice [vzitko2015repulsive, eisenlohr2019mott], $n_{F}(\omega)=(1+e^{\beta\omega})^{-1}$ is the Fermi–Dirac distribution, and $\sigma_{0}$ is a material-dependent constant. This expression holds in the $z\rightarrow\infty$ limit of the Bethe lattice.
@@
The optical conductivity is computed for various values of inter-orbital coupling $t_{\perp}$ and inverse temperature $\beta$, as shown in Fig. 5. For $t_{\perp}=0.7$, we observe three distinct features: (1) a low-frequency Drude peak, originating from coherent quasiparticle excitations near the Fermi level; (2) a mid-infrared peak arising from transitions between the lower Hubbard band and the quasiparticle resonance, and (3) a high-energy charge-excitation peak corresponding to transitions between the lower and upper Hubbard bands. With decreasing temperature, the Drude peak becomes narrower and its height increases, whereas the charge-excitation peak remains largely temperature independent, similar to the earlier reports in the doped infinite-dimensional Hubbard model [jarrell1995optical]. As $t_{\perp}$ approaches the critical value $t_{\perp,c}$, the mid-IR peak is suppressed. For $t_{\perp}>t_{\perp,c}$, an optical gap develops due to the absence of coherent quasiparticles.
@@
In the metallic regime, the optical conductivity spectrum for $\omega\geq 0$ is composed of three distinct features. The first is a Drude peak at zero frequency. The other two features arise from three possible intraband transitions: $\Delta_{ul}$, $\Delta_{um}$, and $\Delta_{ml}$. Of these, the $\Delta_{um}$ and $\Delta_{ml}$ transitions are degenerate, combining to form a single mid-infrared (MIR) peak. The remaining transition, $\Delta_{ul}$, gives rise to the charge excitation peak. Consequently, the total optical spectrum clearly exhibits these three peaks. The charge-excitation peak in $\sigma(\omega)$, corresponding to the transition from the lower to the upper band, is minimum when $\epsilon=0$, and is approximately given by $\Delta_{ul}(\epsilon=0)=t_{\perp}+\sqrt{t_{\perp}^{2}+2ZV^{2}}\approx 2t_{\perp}$. It is interesting to note that, in the limit $t_{\perp}=0$, the MPAM reduces to the PAM, and $\Delta_{ul}\approx\sqrt{2ZV^{2}}$, which is smaller by a factor of $\sqrt{2}$ compared to the gap obtained for the PAM [logan2005dynamics] ($2\sqrt{ZV^{2}}$). This discrepancy arises from the approximation used in determining the roots of the MPAM (see Eq. 10), where the term $\frac{ZV^{2}}{2\omega t_{\perp}}$ is neglected. Including this term would yield the same result as the PAM in the limit $t_{\perp}\rightarrow 0$.
@@
The MIR peak originates from either $\Delta_{um}$ or $\Delta_{ml}$. Their contributions are identical due to particle–hole symmetry, with the minimum excitation occurring at $\epsilon=-\min(D,t_{\perp})$ for $\Delta_{um}$ and at $\epsilon=\min(D,t_{\perp})$ for $\Delta_{ml}$. Hence, the position of the MIR peak, $\omega_{\text{MIR}}$, is proportional to $\Delta_{um}(\epsilon=-\text{min}(D,t_{\perp}))=\Delta_{ml}(\epsilon=\text{min}(D,t_{\perp}))\approx\sqrt{|\text{min}(D,t_{\perp})-t_{\perp}|^{2}+2ZV^{2}}$. As the system approaches the critical point, both $Z\rightarrow 0$ and $|\text{min}(D,t_{\perp})-t_{\perp}|$ decrease. Consequently, the MIR peak redshifts towards $\omega=0$ and the charge-excitation peak blue-shifts away from $\omega=0$ as the system approaches the QCP. In particular, $\omega_{\text{MIR}}$ reaches its minimum at the QCP with an exponent of $0.62$, as shown in the top panel of Fig. 6. Notably, the MIR peak persist even in the $Z\rightarrow 0$ limit due to the finite contribution from $|D-t_{\perp,c}|$ ($t_{\perp,c}=1.033>D$). This trend is evident in Fig. 7, where the MIR peak moves towards $\omega=0$ as $t_{\perp}$ increases, reaching a minimum at $t_{\perp}=t_{\perp,c}$. For $t_{\perp}<D=1.0$, however, $\omega_{\text{MIR}}\propto\sqrt{2ZV^{2}}$, which is explicitly independent of $t_{\perp}$. This implies that the MIR peak provides information about the hybridization gap and quasi-particle weight, similar to the case in the PAM [vidhyadhiraja2005optical]. However, as $t_{\perp}\rightarrow t_{\perp,c}$, other model parameters (here, $t_{\perp}$) can also influence the MIR peak. Beyond the QCP ($t_{\perp}>t_{\perp c}$), the MIR peak disappears. At the QCP, due to the diverging nature of $\Sigma(\omega)$, the spectral functions, $A(\omega,\epsilon)\propto\omega^{r}$, with $r=0.36$ [kksujan2023emergent] being the soft-gap exponent. This results in a power-law behavior of the optical conductivity with an exponent $2r$. The bottom panel of Fig. 5 shows that as the temperature is decreased, the optical conductivity approaches a power-law form with an exponent $0.737$, which is approximately $2r$.
@@
Upon examining the top panel of Fig. 5, we identify two crossing points, commonly referred to as isosbestic points (ISPs) [greger2013isosbestic]. In the present context, these correspond to temperature-independent crossings in the optical conductivity for a given set of parameters. Similar isosbestic behavior has been reported previously [greger2013isosbestic, jarrell1995optical]. In the regime $t_{\perp}<t_{\perp,c}$, two ISPs are clearly visible: one located below the mid-infrared (MIR) peak ($\omega_{1}^{\text{ISP}}<\omega_{\text{MIR}}$) and another above it ($\omega_{2}^{\text{ISP}}>\omega_{\text{MIR}}$). These ISPs gradually disappear as the system approaches the QCP. We track them by extracting the corresponding frequencies, $\omega_{1/2}^{\text{ISP}}$. The bottom panel of Fig. 6 presents the evolution of ISPs as a function of $t_{\perp}$. We observe that $\omega_{1}^{\text{ISP}}$ vanishes linearly at the QCP, while $\omega_{2}^{\text{ISP}}$ disappears before reaching the QCP. Observing these distinct behaviors experimentally would offer compelling evidence for the existence of a QCP.
Our analysis of transport properties reveals clear signatures of Mott quantum criticality in the optical conductivity. We now turn to the DC resistivity to investigate its quantum critical scaling behavior.
@@ III.3.2 DC conductivity and resistivity scaling
Figure 8: The DC conductivity, $\sigma^{\text{DC}}(T)$, computed using Eq. 8, is plotted for several values of the inter-orbital coupling $t_{\perp}$. The blue dashed line denotes the result from a toy model calculation at the QCP, obtained by substituting the power-law form of the self-energy, $\Sigma_{\text{f}}(\omega)\propto T^{-r}|\omega/T|^{-r}$, into Eq. 8. As $T\rightarrow 0$ at the QCP, the toy model DC conductivity matches $\sigma^{\text{DC}}(T,t_{\perp,c})$. Parameters: $U=1.75$.
The DC conductivity is determined by taking the limit as $\omega\rightarrow 0$ in the optical conductivity expression ( 7), which then reduces to the following form:
$\frac{\sigma^{\text{DC}}(T)}{\sigma_{0}}$ $=2\pi\int d\epsilon\,\text{A}_{0}(\epsilon)\int d\omega^{\prime}\left(-\frac{dn_{F}}{d\omega^{\prime}}\right)$
@@
DC conductivity, $\sigma^{\text{DC}}(T)$, is plotted for different inter-orbital couplings, $t_{\perp}$, in Fig. 8. As expected, $\sigma^{\text{DC}}(T)$ diverges for $t_{\perp}<t_{\perp,c}$ and vanishes for $t_{\perp}>t_{\perp,c}$. At the QCP, it follows a different power-law behavior. At the QCP, the characteristic $\omega/T$ scaling implies a power-law form for the self-energy, $\Sigma_{\text{f}}(\omega)\propto T^{-r}|\omega/T|^{-r}$. This, in turn, governs the behavior of the spectral functions, resulting in the relation $\text{A}_{\text{c}}(\omega,\epsilon)=\text{A}_{\text{c}_{M}}(\omega,\epsilon)\propto T^{2r}$. This leads to the DC conductivity scaling as $\sigma_{\text{DC}}\propto T^{2r}$ at the QCP. A crude estimate of the DC conductivity can thus be obtained from Eq. 8 by assuming a toy model self-energy form, $\Sigma_{\text{f}}(\omega)\propto T^{-r}|\omega/T|^{-r}$. Such a toy model calculation suggests that very low temperatures, as low as $T\sim 10^{-5}$, are required to observe this scaling clearly. However, such calculations suffer from oscillations at very low frequencies and temperatures [vzitko2015repulsive]. Ideally, accurate results would be obtained in the limit $\Lambda\rightarrow 1$ with a larger $n_{\text{states}}$, but this remains infeasible with current computational resources.
Figure 9: (Main) Scaled DC resistivity, $\rho^{DC}(T)/\rho^{DC}_{c}(T)$ is plotted as a function of scaled temperature($T/T_{0}$) for different values of $U$. (Inset) $T_{0}$ is plotted as a function of $\delta U=|U-U_{C}|$, where $U_{c}=1.75$. $T_{0,\text{M}},T_{0,\text{I}}$ corresponds to the scaled temperatures $T_{0}$ in Metal and Mott insulating side. $\text{T}_{0,\text{M}},T_{0,\text{I}}$ vanishes with exponents ($0.76,0.66$). Parameters used: $t_{\perp}=t_{\perp\text{c}}=1.033$.
Resistivity is obtained from the inverse of the DC conductivity, $\rho^{\text{DC}}=1/\sigma^{\text{DC}}$. Systems exhibiting Mott quantum criticality display a characteristic scaling behavior in the electrical resistivity [terletska2011quantum, vuvcivcevic2013finite, eisenlohr2019mott]. This scaling provides insights into the quantum critical region and allows for the extraction of the critical exponent $\nu z$, where $\nu$ is the correlation length exponent and $z$ is the dynamical exponent. A hallmark of Mott quantum criticality is the scaling collapse of the DC resistivity curves. This universal behavior is captured by the scaling form: